how to simplify values like tan225

Grade:

3 Answers

vikhyat chaudhry
12 years ago

tan225= tan(270-45) = cot 45 = 1

Godfrey Classic Prince
12 years ago

Dear Sameer Chowdary,

This can be simplified like so

tan2250= tan(270-45) = cot 450 = 1

(or)

tan2250=tan(180+45)=tan(180+45)=tan 450=1

Hope this helped you immmensely...

Please approve my answer if you liked it by clicking on "Yes" given below...!!Smile

Rohan Das
10 years ago

you can solve this in two ways.

one; tan(225)=tan(180+45)=tan45=1

two  tan(225)=tan(270-45)=cot45=1